# Latent Multimodal Reconstruction for Misinformation Detection

### v1

@@ 8 Conclusions
In this study, we address the challenge of MMD by introducing the “MisCaption This!” dataset, and the Latent Multimodal Reconstruction (LAMAR) method. To create “MisCaption This!”, we use an LVLM to generate diverse and realistic synthetic miscaptioned images, enhancing training data quality and by extension model generalization. LAMAR employs a reconstruction-based approach, where the original truthful caption embeddings are reconstructed from manipulated image-caption pairs, providing an auxiliary signal to aid the detection process. Extensive experiments show that models trained on “MisCaption This!” achieve superior out-of-distribution generalization on the real-world VERITE benchmark, outperforming named entity manipulations and cross-modal misalignment by 7.8% and 10.4%, respectively. Additionally, LAMAR set new state-of-the-art performance on all three VERITE tasks—“True vs. OOC”, “True vs. MC”, and multiclass classification—outperforming prior SotA by 4.3%, 3.0%, and 4.8%, respectively. These results highlight the potential usefulness of leveraging LVLMs for dataset generation and the potential of reconstruction networks in improving MMD performance.
Our study represents a first step toward leveraging LVLMs to enhance training data quality for MMD but is limited to leveraging a single model, LLaVa 1.6 7B. Future research could explore alternative LVLMs, jailbreak prompts, prompt tuning or few-shot prompting to further improve the robustness of generated training data. Nevertheless, researchers should consider potential risks, particularly in inadvertently enabling malicious actors to refine AI-generated disinformation campaigns. Furthermore, future research could further refine the reconstruction network by integrating additional external information or evidence to enhance the detection process. We do not explore this avenue due to the absence of datasets containing external evidence for miscaptioned images. A key challenge in collecting and leveraging external information is the risk of information leakage, which can undermine the early detection of emerging misinformation [42, 31]. Additionally, existing evidence-based OOC datasets often contain artifacts that models may exploit as shortcuts rather than truly assessing factual consistency [43]. Addressing these limitations would be a valuable step forward for the fields of MMD and automated fact-checking.
@@ Acknowledgments
This work is partially funded by the project “vera.ai: VERification Assisted by Artificial Intelligence” under grant agreement no. 101070093 and “DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies” under grant agreement no. 101079164.

### v3

@@ 9 Conclusion
In this study, we address the challenge of Multimodal Misinformation Detection (MMD) by introducing the “MisCaption This!” framework and the Latent Multimodal Reconstruction (LAMAR) architecture. “MisCaption This!” leverages Vision-Language Models (VLMs) to generate diverse, realistic synthetic miscaptioned images, significantly improving training data quality and ultimately model generalization. The LAMAR architecture adopts a reconstruction-based approach: it learns to reconstruct the embedding of the original, truthful caption from manipulated image-caption pairs, thereby offering an effective auxiliary signal that enhances detection. Extensive experiments demonstrate that models trained on “MisCaption This!” data significantly outperform previous methods – including those based on named entity manipulation and cross-modal misalignment – by 7.8% and 10.4% on the VERITE benchmark, respectively. Furthermore, LAMAR establishes new SotA performance across all three VERITE tasks: True vs. OOC (+4.3%), True vs. MC (+3.0%), and multiclass classification (+5.6%).
Despite the promising gains achieved by “MisCaption This!” and LAMAR, this work stresses that MMD remains an open and challenging task. Apart from high “out-of-context” performance ($\approx 85\%$), our final model achieved $\approx 66\%$ and $\approx 55\%$ in the miscaptioned and multiclass scenarios, respectively, which indicates considerable room for improvement. We argue that while assessing the internal consistency of the image-caption pair is a crucial prerequisite for MMD, it is rarely sufficient. Therefore, a critical future direction involves collecting and integrating external knowledge and evidence (e.g., related news articles) for comprehensive fact-checking. Although evidence-enriched datasets exist, such as the enriched NewsCLIPpings dataset for out-of-context cases [1], they often suffer from significant issues including information leakage, dataset-specific biases [29], or unreliable evidence [10, 9]. This emphasizes the pressing need for a new, robust, evidence-enriched dataset specifically targeting miscaptioned images to accurately model and address external inconsistency. Synthetic data generation through “MisCaption This!”, and reconstruction-based architectures, such as LAMAR, could play an instrumental role in advancing these research directions.
@@ Acknowledgments
This work is partially funded by the projects “vera.ai: VERification Assisted by Artificial Intelligence” under grant agreement no. 101070093, “DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies” under grant agreement no. 101079164, and “AI4TRUST - AI-based-technologies for trustworthy solutions against disinformation” under grant agreement no. 101070190.